QUESTION 1
Solve the following simultaneous equations using inverse method:
2x – 3y + z = -2
x - 6y +3z = -2
3x+ 3y - 2z = 2

QUESTION 2
The intersection point between two straight lines is (3,0), where the slope is given as below:
M1 = -2 dan M2 = 1
Find the equation for each of the two straight lines.

Answer by madhan_math(34)

QUESTION 2
The intersection point between two straight lines is (3,0), where the slope is given as below:
M1 = -2 dan M2 = 1
Find the equation for each of the two straight lines?
Solution:
We know that the general equation of straight line is y=mx+c , where 'm' is the slope and c is the constant.

Here given two slopes of the straight lines.
i.e., m1=-2 and m2=1
Therefore,
y = -2x+c1 -----(1)
and y = x+c2 ------(2)
Now we have to find constant terms. That is the reason they gave intersection point.
We use the intersection point (3,0) to find the two constants.
First we substitute the point in equation (1)
0=-2(3)+c1 implies -6+c1=0
Therefore c1=6-------(3)
and now equation (2)
0=1+c2 implies c2=-1------(4)

Hence the two straight line equations are
y=-2x+6
and y=x-1 .